# How Transformers Implement Induction Heads: Approximation and Optimization Analysis

### v2

@@ Appendix B Proofs of Optimization Dynamics: Training Stage I
$\begin{pmatrix}x_{s}\\ \sum_{\tau=1}^{s-1}{\rm softmax}_{s}(-\tilde{p}(s-1-\tau))x_{\tau}\end{pmatrix},$
where ${\rm softmax}_{s}(-\tilde{p}(s-1-\tau)):=\frac{\exp(-\tilde{p})(s-1-\tau)}{\sum_{t=2}^{s-1}\exp(-\tilde{p})(s-1-t)}$, and the target function and output of transformer are as follows
$f^{*}(X)=\begin{pmatrix}\frac{\alpha^{\star}}{1+\alpha^{\star}}x_{L-2}\\ \frac{1}{1+\alpha^{\star}}{\rm softmax}({w^{\star}}^{2}x_{L}x_{s-1})x_{s}\end{pmatrix},$
$f_{\theta}(X)$ $=\begin{pmatrix}g(0)\sum_{s=2}^{L-1}{\rm softmax}(-p(0)(L-1-s))\cdot\left(\sum_{\tau=1}^{s-1}{\rm softmax}_{s}(-\tilde{p}(s-1-\tau))x_{\tau}\right)\\ h(0)\sum_{s=2}^{L-2}{\rm softmax}\bigg{(}w_{K}(0)w_{Q}(0)x_{L}\left(\sum_{\tau=1}^{s-1}{\rm softmax}_{s}(-\tilde{p}(s-1-\tau))x_{\tau}\right)\bigg{)}x_{s}\end{pmatrix}$
$=\begin{pmatrix}g(0)\frac{1}{L-2}\sum_{\tau=1}^{L-2}\left(\sum_{s=\tau+1}^{L-1}{\rm softmax}_{s}(-\tilde{p}(s-1-\tau))\right)x_{\tau}\\ h(0)\frac{1}{L-2}\sum_{s=2}^{L-2}x_{s}\end{pmatrix}.$
where ${\rm softmax}:={\rm softmax}_{L}$ is for the second layer.
@@
$\mathcal{L}(\theta)$ $=\mathop{\mathbb{E}}\limits_{X\sim\mathbb{N}(0,1)^{L}}\Bigg{[}\frac{{\alpha^{\star}}^{2}}{(1+\alpha^{\star})^{2}}x_{L-2}^{2}+\frac{g(0)^{2}}{(L-2)^{2}}\sum_{\tau=1}^{L-2}\left(\sum_{s=\tau+1}^{L-1}{\rm softmax}_{s}(-\tilde{p}(s-1-\tau))\right)^{2}x_{\tau}^{2}$
$\quad\quad\quad\quad\quad\quad+\frac{2g(0)}{L-2}\frac{\alpha^{\star}}{1+\alpha^{\star}}{\rm softmax}_{L-1}(0)x_{L-2}^{2}\Bigg{]}+C(w^{\star},\alpha^{\star},w(0),h(0))$
@@
Function I. This function is purely defined for the calculation of $\frac{{\textnormal{d}}p}{{\textnormal{d}}t}$. Denoted by $q(\tilde{p}):=\sum_{\tau=1}^{L-2}\left(\sum_{s=\tau+1}^{L-1}\frac{e^{\tilde{p}(s-1-\tau)}}{\sum_{k=0}^{s-2}e^{-\tilde{p}k}}\right)^{2}$, we first prove $q^{\prime}(\tilde{p})\leq 0$.
@@
$q^{\prime}(\tilde{p})$ $=2(1-e^{-\tilde{p}})e^{-\tilde{p}}\sum_{\tau=1}^{L-2}e^{2\tilde{p}\tau}\left(\sum_{s=\tau+1}^{L-1}\frac{1}{e^{\tilde{p}(s-1)}-1}\right)^{2}$
@@
$q^{\prime}(\tilde{p})$’s last factor can be formed as
@@
where $t=e^{-\tilde{p}}\geq 1$. Since $s\geq\tau+1$, $q^{\prime}(\tilde{p})\leq 0$.
@@
$\frac{{\textnormal{d}}}{{\textnormal{d}}p}{\rm softmax}(-ps)=\frac{{\textnormal{d}}}{{\textnormal{d}}p}\frac{\exp(-ps)}{\sum_{\tau=0}^{L-3}\exp(-p\tau)}=\frac{-s\exp(-ps)M(p)+\exp(-ps)m(p)}{M(p)^{2}}.$
@@
Appendix C Proofs of Optimization Dynamics: Training Stage II
@@
Lemma C.1 (Restate of Lemma 5.4).
@@
$f^{\star}(X)$ $=\begin{pmatrix}\frac{{\alpha^{\star}}}{1+{\alpha^{\star}}}x_{L-2}\\ \frac{1}{1+{\alpha^{\star}}}\sum_{s=2}^{L-1}{\rm softmax}\left({w^{\star}}^{2}x_{L}x_{s-1}\right)x_{s}\end{pmatrix},$
${\textnormal{{TF}}}(X;\theta)$ $=\begin{pmatrix}g\sum_{s=2}^{L-2}{\rm softmax}(-p(L-1-s))x_{s-1}\\ h\sum_{s=2}^{L-2}{\rm softmax}\left(w^{2}x_{L}x_{s-1}\right)x_{s}\end{pmatrix}.$
@@
$=\frac{1}{2}{\mathbb{E}}_{X}\Bigg{[}\left(\frac{{\alpha^{\star}}}{1+{\alpha^{\star}}}x_{L-2}-g\sum_{s=2}^{L-2}{\rm softmax}(-p(L-1-s))x_{s-1}\right)^{2}\Bigg{]}$
$\quad+\frac{1}{2}{\mathbb{E}}_{X}\Bigg{[}\left(\frac{1}{1+{\alpha^{\star}}}\sum_{s=2}^{L-1}{\rm softmax}({w^{\star}}^{2}x_{L}x_{s-1})x_{s}-h\sum_{s=2}^{L-2}{\rm softmax}(w^{2}x_{L}x_{s-1})x_{s}\right)^{2}\Bigg{]}.$
@@
$\mathcal{L}_{\textnormal{{G}}_{4}}(\theta)=\mathcal{L}_{\textnormal{{G}}_{4}}(p,g)=\frac{1}{2}{\mathbb{E}}_{X}\Bigg{[}\left(\frac{{\alpha^{\star}}}{1+{\alpha^{\star}}}x_{L-2}-g\sum_{s=2}^{L-2}{\rm softmax}(-p(L-1-s))x_{s-1}\right)^{2}\Bigg{]},$
$\mathcal{L}_{\textnormal{{IH}}_{2}}(\theta)=\mathcal{L}_{\textnormal{{IH}}_{2}}(w,h)=\frac{1}{2}\mathbb{E}_{X}\Bigg{[}\left(\frac{1}{1+{\alpha^{\star}}}\sum_{s=2}^{L-1}{\rm softmax}({w^{\star}}^{2}x_{L}x_{s-1})x_{s}-h\sum_{s=2}^{L-2}{\rm softmax}(w^{2}x_{L}x_{s-1})x_{s}\right)^{2}\Bigg{]}.$
@@
C.1 Dynamics of the parameters for $4$-gram
@@
Then, a straightforward calculation, combined with Lemma D.3 and Lemma D.4, yields the explicit formulation of $\mathcal{L}_{\textnormal{{G}}_{4}}(\theta)$ and the GF dynamics of $p$ and $g$:
@@
which means $g$ has entered a stable region (although it is possible that $g$ is non-monotonic), while $p$ keeps increase. In fact, if $\tilde{t}:=\inf\{t>0:g(t)=2g^{\star}\}$, then $\frac{{\textnormal{d}}g}{{\textnormal{d}}t}|_{\tilde{t}}<0$, which leads to a contradiction. If $\hat{t}:=\inf\{t>0:\frac{{\textnormal{d}}p(t)}{{\textnormal{d}}t}=0\}$, then
$\left.\left(g^{\star}-g\frac{1+e^{-p(L-2)}}{1+e^{-p}}+g\frac{m(2p)}{m(p)}\right)\right|_{\hat{t}}=0,\quad\frac{{\textnormal{d}}g}{{\textnormal{d}}t}<0,$
@@
C.2 Dynamics of the parameters for induction head
@@
$\mathcal{L}_{\textnormal{{IH}}_{2}}(\theta)=\frac{1}{2}{\mathbb{E}}_{X}\Bigg{[}\left(\frac{1}{1+{\alpha^{\star}}}\sum_{s=2}^{L-1}{\rm softmax}({w^{\star}}^{2}x_{L}x_{s-1})x_{s}-h\sum_{s=2}^{L-2}{\rm softmax}(w^{2}x_{L}x_{s-1})x_{s}\right)^{2}\Bigg{]}.$
Technical simplification. Unlike $\mathcal{L}_{\textnormal{{G}}_{4}}(\theta)$, the denominators of the softmax terms ${\rm softmax}({w^{\star}}^{2}x_{L}x_{s-1})$ and ${\rm softmax}({w}^{2}x_{L}x_{s-1})$ in $\mathcal{L}_{\textnormal{{IH}}_{2}}(\theta)$ depend on the input tokens $X$, making it hard to derive a closed-form expression for $\mathcal{L}_{\textnormal{{IH}}_{2}}(\theta)$. In Bai et al. [2023], the authors consider a simplified transformer model, which replaces the softmax ${\rm softmax}(z_{1},\cdots,z_{L})$ with $\frac{1}{L}\exp(z_{1},\cdots,z_{L})$. This approximation is nearly tight when $z_{1},\cdots,z_{L}\approx 0$. Notice that 1) ${w}^{2}x_{L}x_{s-1}\approx 0$ holds near the small initialization, i.e., for $w\approx\sigma_{\rm init}\ll 1$. In fact, our analysis shows that $w\approx\sigma_{\rm init}$ is maintained over a long period. 2) ${w}^{\star}=\mathcal{O}(1)$, which implies that ${w}^{2}x_{L}x_{s-1}\approx 0$ for most input sequence. Thus, we adopt the simplification used in Bai et al. [2023], resulting in the following approximation of the loss function:
@@
Then by a straightforward calculation with Lemma D.3, we can derive its explicit formulation:
@@
We denote the first hitting time of $h$ decreasing to $h^{\star}$ as $\dot{t}$:
$\dot{t}:=\inf\left\{t>T_{2}^{h}:h(t)<h^{\star}\right\}.$
If $w(\dot{t})\geq w^{\star}$, then at the first hitting time of $w$ increasing to $w^{\star}$, $\frac{{\textnormal{d}}w}{{\textnormal{d}}t}<0$, which leads to a contradiction. If $w(\dot{t})<w^{\star}$, then $\frac{{\textnormal{d}}h}{{\textnormal{d}}t}|_{\dot{t}}>0$, which also leads to a contradiction. Hence, $\dot{t}=+\infty$, which means that $h$ always remains above $h^{\star}$ for $t>T_{2}^{h}$.
@@
Therefore, there exists a small constant $\varepsilon(w(0),w^{\star})$ only depends on $w(0)$ and $w^{\star}$ such that $h$ is dominated by $1+\varepsilon(w(0),w^{\star})$ times right hand side of (24), from which we deduce that (C.2) is a tight estimation of $T_{1}^{h}$ instead of an upper bound, i.e. $T_{1}^{h}=\Theta(L)$.
@@
From the above inequality, (C.2) is not only an upper bound, but a tight estimation of $T_{1/2}^{w}$, i.e.
@@
C.3 Proof of Theorem 5.5
This theorem is a direct corollary of our analysis of the entire training dynamics in Appendix C.1 and C.2, leveraging the relationship between the parameters and the loss.
@@
Recalling our proof in Appendix C.2, for $t<T_{1/2}^{h}=\mathcal{O}(L)$, it holds that $h(t)<\sigma_{{\rm init}}+\mathcal{O}({t}/{((1+\alpha^{\star})L)}),w(t)<\sigma_{{\rm init}}+o({t}/{((1+\alpha^{\star})L)})$. Additionally, since $T_{\rm I}=\Theta(1)\ll\Theta(L)$, it follows that
@@
Appendix D Useful Inequalities
Lemma D.1 (Corollary A.7 in Edelman et al. [2022]).

### v3

@@ D.1 Optimization Dynamics in Training Stage I
$\begin{pmatrix}x_{s}\\ (x_{\tau})_{\tau=1}^{s-1}\ {\rm softmax}\left(\big{(}-\tilde{p}(s-1-\tau)\big{)}_{\tau=1}^{s-1}\right)^{\top}\end{pmatrix},$
and the target function and output of transformer are as follows
$f^{*}(X)=\begin{pmatrix}\frac{\alpha^{\star}}{1+\alpha^{\star}}x_{L-2}\\ \frac{1}{1+\alpha^{\star}}(x_{s})_{s=2}^{L-1}\ {\rm softmax}\Big{(}\big{(}x_{L}{w^{\star}}^{2}x_{s-1}\big{)}_{s=2}^{L-1}\Big{)}^{\top}\end{pmatrix},$
$f_{\theta}(X)$ $=\begin{pmatrix}g(0)\Bigg{(}(x_{\tau})_{\tau=1}^{s-1}{\rm softmax}\left((-\tilde{p}(s-1-\tau))_{\tau=1}^{s-1}\right)^{\top}\Bigg{)}_{s=2}^{L-1}{\rm softmax}\Big{(}(-p(0)(L-1-s))_{s=2}^{L-1}\Big{)}^{\top}\\ h(0)(x_{s})_{s=2}^{L-2}{\rm softmax}\Bigg{(}\Big{(}w_{K}(0)w_{Q}(0)x_{L}\cdot(x_{\tau})_{\tau=1}^{s-1}\ {\rm softmax}\left((-\tilde{p}(s-1-\tau))_{\tau=1}^{s-1}\right)^{\top}\Big{)}_{s=2}^{L-2}\Bigg{)}^{\top}\end{pmatrix}$
$=\begin{pmatrix}g(0)\frac{1}{L-2}\sum_{\tau=1}^{L-2}\left(\sum_{s=\tau+1}^{L-1}{\rm softmax}\Big{(}(-\tilde{p}(s-1-t))_{t=1}^{s-1}\Big{)}\right)_{t=\tau}x_{\tau}\\ h(0)\frac{1}{L-2}\sum_{s=2}^{L-2}x_{s}\end{pmatrix}.$
@@
$\mathcal{L}(\theta)$ $=\mathop{\mathbb{E}}\limits_{X\sim\mathbb{N}(0,1)^{L}}\Bigg{[}\frac{{\alpha^{\star}}^{2}}{(1+\alpha^{\star})^{2}}x_{L-2}^{2}+\frac{g(0)^{2}}{(L-2)^{2}}\sum_{\tau=1}^{L-2}\left(\sum_{s=\tau+1}^{L-1}{\rm softmax}\Big{(}(-\tilde{p}(s-1-t))_{t=1}^{s-1}\Big{)}_{t=\tau}\right)^{2}x_{\tau}^{2}$
$\quad\quad\quad\quad\quad\quad+\frac{2g(0)}{L-2}\frac{\alpha^{\star}}{1+\alpha^{\star}}{\rm softmax}\Big{(}(-p(0)(L-1-s))_{s=2}^{L-1}\Big{)}_{s=L-1}x_{L-2}^{2}\Bigg{]}+C(w^{\star},\alpha^{\star},w(0),h(0))$
@@
Function I. This function is purely defined for the calculation of $\frac{{\textnormal{d}}p}{{\textnormal{d}}t}$. Denoted by $q(\tilde{p}):=\sum_{\tau=1}^{L-2}\left(\sum_{s=\tau+1}^{L-1}\frac{e^{\tilde{p}(s-1-\tau)}}{\sum_{k=0}^{s-2}e^{-\tilde{p}k}}\right)^{2}$, we first prove $\frac{{\textnormal{d}}q}{{\textnormal{d}}\tilde{p}}\leq 0$.
@@
$\frac{{\textnormal{d}}q}{{\textnormal{d}}\tilde{p}}$ $=2(1-e^{-\tilde{p}})e^{-\tilde{p}}\sum_{\tau=1}^{L-2}e^{2\tilde{p}\tau}\left(\sum_{s=\tau+1}^{L-1}\frac{1}{e^{\tilde{p}(s-1)}-1}\right)^{2}$
@@
$\frac{{\textnormal{d}}q}{{\textnormal{d}}\tilde{p}}$’s last factor can be formed as
@@
where $t=e^{-\tilde{p}}\geq 1$. Since $s\geq\tau+1$, $\frac{{\textnormal{d}}q}{{\textnormal{d}}\tilde{p}}\leq 0$.
@@
$\frac{{\textnormal{d}}}{{\textnormal{d}}p}{\rm softmax}\Big{(}(-p(L-1-t))_{t=2}^{L-1}\Big{)}_{t=L-1-s}=\frac{{\textnormal{d}}}{{\textnormal{d}}p}\frac{\exp(-ps)}{\sum_{\tau=0}^{L-3}\exp(-p\tau)}=\frac{-s\exp(-ps)M(p)+\exp(-ps)m(p)}{M(p)^{2}}.$
@@
D.2 Optimization Dynamics in Training Stage II
@@
Lemma D.1 (Restate of Lemma 4.4).
@@
$f^{\star}(X)$ $=\begin{pmatrix}\frac{{\alpha^{\star}}}{1+{\alpha^{\star}}}x_{L-2}\\ \frac{1}{1+{\alpha^{\star}}}(x_{s})_{s=2}^{L-1}{\rm softmax}\Big{(}\left({w^{\star}}^{2}x_{L}x_{s-1}\right)_{s=2}^{L-1}\Big{)}^{\top}\end{pmatrix},$
${\textnormal{{TF}}}(X;\theta)$ $=\begin{pmatrix}g\cdot(x_{s-1})_{s=2}^{L-1}{\rm softmax}\Big{(}(-p(L-1-s))_{s=2}^{L-1}\Big{)}^{\top}\\ h\cdot(x_{s})_{s=2}^{L-1}{\rm softmax}\Big{(}\left(w^{2}x_{L}x_{s-1}\right)_{s=2}^{L-1}\Big{)}^{\top}\end{pmatrix}.$
@@
$=\frac{1}{2}{\mathbb{E}}_{X}\Bigg{[}\left(\frac{{\alpha^{\star}}}{1+{\alpha^{\star}}}x_{L-2}-g\cdot(x_{s-1})_{s=2}^{L-1}{\rm softmax}\Big{(}(-p(L-1-s))_{s=2}^{L-1}\Big{)}^{\top}\right)^{2}\Bigg{]}$
$\quad+\frac{1}{2}{\mathbb{E}}_{X}\Bigg{[}\left(\frac{1}{1+{\alpha^{\star}}}(x_{s})_{s=2}^{L-1}{\rm softmax}\Big{(}\left({w^{\star}}^{2}x_{L}x_{s-1}\right)_{s=2}^{L-1}\Big{)}^{\top}-h\cdot(x_{s})_{s=2}^{L-1}{\rm softmax}\Big{(}\left(w^{2}x_{L}x_{s-1}\right)_{s=2}^{L-1}\Big{)}^{\top}\right)^{2}\Bigg{]}.$
@@
$\mathcal{L}_{\textnormal{{G}}_{4}}(\theta)$ $=\mathcal{L}_{\textnormal{{G}}_{4}}(p,g)$
$=\frac{1}{2}{\mathbb{E}}_{X}\Bigg{[}\left(\frac{{\alpha^{\star}}}{1+{\alpha^{\star}}}x_{L-2}-g\cdot(x_{s-1})_{s=2}^{L-1}{\rm softmax}\Big{(}(-p(L-1-s))_{s=2}^{L-1}\Big{)}\right)^{\top})^{2}\Bigg{]},$
$\mathcal{L}_{\textnormal{{IH}}_{2}}(\theta)$ $=\mathcal{L}_{\textnormal{{IH}}_{2}}(w,h)$
$=\frac{1}{2}\mathbb{E}_{X}\Bigg{[}\left(\frac{1}{1+{\alpha^{\star}}}(x_{s})_{s=2}^{L-1}{\rm softmax}\Big{(}\left({w^{\star}}^{2}x_{L}x_{s-1}\right)_{s=2}^{L-1}\Big{)}^{\top}-h\cdot(x_{s})_{s=2}^{L-2}{\rm softmax}\Big{(}\left(w^{2}x_{L}x_{s-1}\right)_{s=2}^{L-1}\Big{)}^{\top}\right)^{2}\Bigg{]}.$
@@
D.2.1 Dynamics of the parameters for $4$-gram
@@
Then, a straightforward calculation, combined with Lemma E.3 and Lemma E.4, yields the explicit formulation of $\mathcal{L}_{\textnormal{{G}}_{4}}(\theta)$ and the GF dynamics of $p$ and $g$:
@@
which means $g$ has entered a stable region (although it is possible that $g$ is non-monotonic), while $p$ keeps increase. In fact, if $T^{g}_{2g^{\star}}:=\inf\{t>0:g(t)=2g^{\star}\}$, then $\frac{{\textnormal{d}}g}{{\textnormal{d}}t}|_{T^{g}_{2g^{\star}}}<0$, which leads to a contradiction. If $T^{{{\textnormal{d}}p}/{{\textnormal{d}}t}}_{0}:=\inf\{t>0:\frac{{\textnormal{d}}p(t)}{{\textnormal{d}}t}=0\}$, then
$\left.\left(g^{\star}-g\frac{1+e^{-p(L-2)}}{1+e^{-p}}+g\frac{m(2p)}{m(p)}\right)\right|_{T^{{{\textnormal{d}}p}/{{\textnormal{d}}t}}_{0}}=0,\quad\frac{{\textnormal{d}}g}{{\textnormal{d}}t}<0,$
@@
D.2.2 Dynamics of the parameters for induction head
@@
$\mathcal{L}_{\textnormal{{IH}}_{2}}(\theta)={\frac{1}{2}\mathbb{E}_{X}\Bigg{[}\left(\frac{1}{1+{\alpha^{\star}}}(x_{s})_{s=2}^{L-1}{\rm softmax}\Big{(}\left({w^{\star}}^{2}x_{L}x_{s-1}\right)_{s=2}^{L-1}\Big{)}^{\top}-h\cdot(x_{s})_{s=2}^{L-2}{\rm softmax}\Big{(}\left(w^{2}x_{L}x_{s-1}\right)_{s=2}^{L-2}\Big{)}^{\top}\right)^{2}\Bigg{]}}.$
Technical simplification. Unlike $\mathcal{L}_{\textnormal{{G}}_{4}}(\theta)$, the denominators of the softmax terms ${\rm softmax}\Big{(}\left({w^{\star}}^{2}x_{L}x_{s-1}\right)_{s=2}^{L-1}\Big{)}$ and ${\rm softmax}\Big{(}\left(w^{2}x_{L}x_{s-1}\right)_{s=2}^{L-2}\Big{)}$ in $\mathcal{L}_{\textnormal{{IH}}_{2}}(\theta)$ depend on the input tokens $X$, making it hard to derive a closed-form expression for $\mathcal{L}_{\textnormal{{IH}}_{2}}(\theta)$. In Bai et al. (2023), the authors consider a simplified transformer model, which replaces ${\rm softmax}(z_{1},\cdots,z_{L})$ with $\frac{1}{L}\exp(z_{1},\cdots,z_{L})$. This approximation is nearly tight when $z_{1},\cdots,z_{L}\approx 0$. Notice that 1) ${w}^{2}x_{L}x_{s-1}\approx 0$ holds near the small initialization, i.e., for $w\approx\sigma_{\rm init}\ll 1$. In fact, our analysis shows that $w\approx\sigma_{\rm init}$ is maintained over a long period. 2) ${w}^{\star}=\mathcal{O}(1)$, which implies that ${w}^{2}x_{L}x_{s-1}\approx 0$ for most input sequence. Thus, we adopt the simplification used in Bai et al. (2023), resulting in the following approximation of the loss function:
@@
Then by a straightforward calculation with Lemma E.3, we can derive its explicit formulation:
@@
We denote the first hitting time of $h$ decreasing to $h^{\star}$ as $T^{h}_{h^{\star}}$:
$T^{h}_{h^{\star}}:=\inf\left\{t>T_{2}^{h}:h(t)<h^{\star}\right\}.$
If $w(T^{h}_{h^{\star}})\geq w^{\star}$, then at the first hitting time of $w$ increasing to $w^{\star}$, $\frac{{\textnormal{d}}w}{{\textnormal{d}}t}<0$, which leads to a contradiction. If $w(T^{h}_{h^{\star}})<w^{\star}$, then $\frac{{\textnormal{d}}h}{{\textnormal{d}}t}|_{T^{h}_{h^{\star}}}>0$, which also leads to a contradiction. Hence, $T^{h}_{h^{\star}}=+\infty$, which means that $h$ always remains above $h^{\star}$ for $t>T_{2}^{h}$.
@@
Therefore, there exists a small constant $\varepsilon(w(0),w^{\star})$ only depends on $w(0)$ and $w^{\star}$ such that $h$ is dominated by $1+\varepsilon(w(0),w^{\star})$ times right hand side of (24), from which we deduce that (D.2.2) is a tight estimation of $T_{1}^{h}$ instead of an upper bound, i.e. $T_{1}^{h}=\Theta(L)$.
@@
From the above inequality, (D.2.2) is not only an upper bound, but a tight estimation of $T_{1/2}^{w}$, i.e.
@@
D.3 Proof of Theorem 4.5
This theorem is a direct corollary of our analysis of the entire training dynamics in Appendix D.2.1 and D.2.2, leveraging the relationship between the parameters and the loss.
@@
Recalling our proof in Appendix D.2.2, for $t<T_{1/2}^{h}=\mathcal{O}(L)$, it holds that $h(t)<\sigma_{{\rm init}}+\mathcal{O}({t}/{((1+\alpha^{\star})L)}),w(t)<\sigma_{{\rm init}}+o({t}/{((1+\alpha^{\star})L)})$. Additionally, since $T_{\rm I}=\Theta(1)\ll\Theta(L)$, it follows that
@@
Appendix E Useful Inequalities
Lemma E.1 (Corollary A.7 in Edelman et al. (2022)).